when 12 grams of zinc oxide react with 12 grams of hydrochloric acid according to the following balanced chemical equation, a) what is the limiting reactant? b) if 13g of zinc chloride are formed, what is the percent yield of the reaction?

ZnO + 2HCl ------> ZnCl2 + H2O
Answer #1

ZnO + 2HCl = ZnCl2 + H2O

moles of ZnO = mass / molar mass

= 12 / 81.38 = 0.147

moles of HCl = 12 / 36.5 = 0.329

according to reaction

moles of ZnO = 1/2 moles of 2HCl

so, 0.329 /2 = 0.1645

we should have 0.1645 moles of ZnO if we have 12 g of HCL

so, ZnO is limiting reagent and product will form according to the amount of limiting reagent

moles of ZnO = moles of ZnCl2 ( balance reaction ) = 0.1645

mass of ZnCl2 = mole * molar mass

= 0.1645 * 136.3

= 20.1 grams

percent yield =( observed mass / caculated mass )*100

= (13 / 20.1)*100

= 64.68 %

any query please comment

if satisfied please rate it thanks